1. The number of emails that I get in one hour has such density function:
x 2 1 4 5 3 6 and more 0
f(x) 0.17 0.23 0.08 0.04 0.14 0.02 ?
a) find the probability, that there will be no letter for the next hour;
b) find cumulative distributional function.
1
Expert's answer
2020-04-09T15:14:19-0400

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c}
 x & 2 & 1 & 4 & 5 & 3 & 6\text{ and more} & 0 \\ \hline
 f(x) & 0.17 & 0.23 & 0.08 & 0.04 & 0.14 & 0.02 & ? \\
 
\end{array}"


a) "P(X=0)" - the probability, that there will be no letter for the next hour.

To find "P(X=0)" we can use the fact that "\sum P(X=x) =1" .


We have: "0.17+0.23+0.08+0.04+0.14+0.02+P(X=0)=1"

"P( X=0)= 0.32"

Answer: 0.32


b) "F(x)=P(X\leq x)=P(X=0)+P(X=1)+…+P(X=x)"

Using this formula, we have:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c}
 x & 0 & 1 & 2 & 3 & 4 &5& 6\text{ and more} \\ \hline
 F(x) & 0.32 & 0.55 & 0.72 & 0.86 & 0.94 & 0.98 & 1 \\
 
\end{array}"